# Super LiDAR Reflectance for Robotic Perception

### v1

@@ Abstract
Conventionally, human intuition often defines vision as a modality of passive optical sensing, while active optical sensing is typically regarded as measuring rather than the default modality of vision. However, the situation now changes: sensor technologies and data-driven paradigms empower active optical sensing to redefine the boundaries of vision, ushering in a new era of active vision. Light Detection and Ranging (LiDAR) sensors capture reflectance from object surfaces, which remains invariant under varying illumination conditions, showcasing significant potential in robotic perception tasks such as detection, recognition, segmentation, and Simultaneous Localization and Mapping (SLAM). These applications often rely on dense sensing capabilities, typically achieved by high-resolution, expensive LiDAR sensors. A key challenge with low-cost LiDARs lies in the sparsity of scan data, which limits their broader application. To address this limitation, this work introduces an innovative framework for generating dense LiDAR reflectance images from sparse data, leveraging the unique attributes of non-repeating scanning LiDAR (NRS-LiDAR). We tackle critical challenges, including reflectance calibration and the transition from static to dynamic scene domains, facilitating the reconstruction of dense reflectance images in real-world settings. The key contributions of this work include a comprehensive dataset for LiDAR reflectance image densification, a densification network tailored for NRS-LiDAR, and diverse applications such as loop closure and traffic lane detection using the generated dense reflectance images. Experimental results validate the efficacy of the proposed approach, which successfully integrates computer vision techniques with LiDAR data processing, enhancing the applicability of low-cost LiDAR systems and establishing a novel paradigm for robotic active vision—LiDAR as a Camera. The dataset and code are available at: To Be Updated.
@@ I Introduction
Recent advancements in LiDAR technology have greatly enhanced its accessibility, driven by reductions in sensor size and cost, thereby expanding its applicability across a wide range of robotic applications. In addition to delivering precise 3D spatial data, LiDAR systems are capable of capturing reflectance intensity from objects’ surfaces, acquiring illumination-invariant environmental information, and rich texture to complement the limitations of pure geometric-based LiDAR sensing.
@@
Figure 1: Comparison between the reflectance image densified by the proposed method and the true reflectance images stitched from high-line LiDAR data. (a) Raw NRS-LiDAR input. (b) Densified 1380 $\times$ 240 reflectance image using the proposed method. (c) Stitched 1024 $\times$ 128 reflectance images using the high-cost Ouster 128-line LiDAR data from [1, 2]. Thanks to the proposed network and NRS-LiDARs, the method effectively eliminates line artifacts and mitigates overexposure in challenging scenarios, generating high-definition, camera-grade reflectance images.
@@
To date, numerous studies have used LiDAR reflectance intensity for a variety of applications, including loop closure detection [2, 3, 4], intensity-enhanced LiDAR odometry [1, 5, 6, 7], segmentation [8, 9], and object detection [10, 11]. These applications typically depend on dense reflectance images, which are often generated by high-resolution, costly LiDAR sensors such as the Ouster OS1-128 and Velodyne VLS-128. In contrast, low-cost LiDAR sensors, such as the Velodyne-16, Velodyne-32, or Livox MID-360, generally produce sparse data, posing significant challenges for directly applying methods designed for dense reflectance images.
@@
Intuitively, generating dense and accurate reflectance images from sparse LiDAR data holds significant potential to advance robotic perception tasks, yet this area remains underexplored in existing research [12]. Successfully tackling this challenge would effectively bridge the gap between computer vision techniques and LiDAR data, facilitating the adaptation of camera-based methodologies to LiDAR systems. Moreover, it would address the inherent limitations of traditional cameras, such as their sensitivity to lighting conditions, while simultaneously enhancing the texture representation capabilities of LiDAR technology.
@@
Creating dense LiDAR reflectance images typically depends on high-quality, high-resolution ground-truth data, particularly when using deep neural networks. However, acquiring such data often requires specialized sensors or configurations, posing significant challenges. In this study, we capitalize on the distinctive imaging properties of NRS-LiDAR to produce high-quality ground-truth reflectance images (Fig. 1). As shown in Fig. 2, dense ground-truth data can be acquired in static environments through scan accumulation. Nevertheless, acquiring dense ground-truth data is not feasible when the robot is in motion. In dynamic scenarios, sparse and rapidly changing inputs must be processed in real time, significantly increasing the complexity of dense reconstruction. The disparity between static ground-truth generation and dynamic scene processing underscores the challenges of achieving dense reflectance images with low-cost LiDARs. In summary, we face two challenging issues. The first is to acquire dense ground-truth for reflectance image densification in motion conditions. The second issue is how to densify the sparse reflectance given ground-truth data. To address these issues, we propose a novel method that includes the following contributions:
@@
    We propose the first benchmark dataset tailored for LiDAR reflectance image densification. Compare to existing LiDAR datasets [13, 1], our dataset providing paired sparse-to-dense ground truth data, enabling both end-to-end deep learning based densification benchmarking and image-level perception tasks.
@@
    We introduce a novel convolutional network specifically designed for real-time densification of the reflectance image of NRS-LiDAR (e.g., Livox MID-360). Our approach effectively addresses two critical challenges: the static-to-motion domain gap and adaptive intensity calibration in LiDAR reflectance image densification. Compared to existing methods, our method achieves significant improvements in quality and accuracy.
@@
    We validate the effectiveness of the densified reflectance image in loop closure and traffic lane detection, achieving significant improvement under challenging conditions and demonstrating its potential to enhance practical applications of low-cost LiDAR systems.
@@
Figure 2: Illustration of the non-repeating scanning mechanism. When stationary, the LiDAR gradually produces denser point clouds over time within the same FoV, allowing the reflectance intensity image derived from the point cloud projection to achieve higher density as well. The color of point clouds or gray-scale images represents the intensity value.
II Related Works
II-A LiDAR Reflectance in Robotics
@@
In single LiDAR-based applications, numerous studies have primarily focused on exploiting geometric information for robotic perception[14] and localization[15] tasks. Nevertheless, these approaches exhibit limitations in geometrically degenerated environments due to their inability to capture texture information. To address this, researchers have explored fusing camera and LiDAR data to exploit their complementary strengths [16, 17]. However, this integration presents challenges, including additional hardware requirements, precise calibration, and temporal synchronization. Moreover, visual sensors are inherently limited by illumination conditions.
@@
A promising alternative focuses on leveraging LiDAR’s intrinsic reflectance properties. Significant progress has been demonstrated in this domain. A loop closure detection method that synergistically combines geometric and intensity data is proposed [18]. Furthermore, the application of reflectance information has proven particularly effective in robotic terrain analysis, enabling semantic segmentation to distinguish between road surfaces and grasslands[19].
@@
Recently, dense reflectance images, which are generated by high-resolution LiDAR sensors, have been utilized to enhance robotic perception capabilities[5, 6]. A notable work is COIN-LIO[1], which is a LiDAR-inertial odometry framework integrating photometric error minimization on LiDAR intensity images with geometric registration to address challenges in geometrically degenerated environments. A hybrid loop closure detection method is introduced in [2] by combining LiDAR point cloud and LiDAR intensity image, achieving superior robustness. However, these advanced methods rely on high-resolution, high-cost LiDAR sensors, making them impractical for low-cost applications and emphasizing the need for alternative strategies in resource-constrained scenarios.
This study presents a novel approach of reconstructing high-fidelity dense reflectance images from sparse ones with low-cost LiDAR sensors, and demonstrates significant potential in real-world real-time robot applications.
II-B Depth Completion
@@
Dense environmental sensing is critical for robotic tasks. However, physical sensors typically provide only sparse measurements. Completing sparse sensory data into dense perceptual representations is thus a crucial research direction. Depth completion[20, 21] is one of the most extensively studied domains, inspiring us to review dense completion methods. Depth completion utilizes sparse depth data from sensor measurements[22, 23, 24, 25], multi-view geometry[26, 27], or data-driven priors[28, 29], to generate dense depth images, enabling continuous and dense perception for navigation and manipulation.
@@
However, directly applying existing depth completion approaches to LiDAR reflectance images densification is challenging due to the complexities of intensity measurements. Factors such as incidence angle, surface reflectance properties, and sensor-to-target distance[30] complicate the learning of latent intensity patterns, especially with limited public training data available. Therefore, developing novel networks specifically for LiDAR reflectance images densification is essential. Unlike depth completion, which focuses solely on geometry, LiDAR reflectance images densification equips robots with a new, high-resolution eye, thus allowing high-level visual perception to be achieved through LiDAR, opening the door to reimagining various vision problems.
@@ III Methodology
Figure 3: System diagram: During stationary, the robot utilizes NRS-LiDAR to obtain sparse and dense LiDAR data. Sparse data is augmented to simulate motion characteristics, while dense data serves as supervision to train a reflectance image densification network for motion perception tasks.
Robots in motion often require sufficient time to accumulate data for a comprehensive understanding of their environment. In this section, we detail our proposed method to address this challenge. As shown in Fig. 3, the purpose of this work is to rely solely on sparse LiDAR data collected by the robot over a short period during motion as input to generate dense and accurate reflectance images as output, providing comprehensive environmental information for robotic perception tasks.
@@ III-A1 Data Acquisition
The Super LiDAR Reflectance dataset was acquired using a Giraffe ground robot platform equipped with a multi-sensor suite: an Intel RealSense D435i camera capturing RGB images at 1280 × 720 resolution, a Livox MID-360 LiDAR sensor with a 360° horizontal and 59° vertical field of view (FoV) operating at 10 Hz point cloud generation rate, and a high-frequency IMU sampling at 200 Hz. Initially, the robot autonomously navigated through selected representative environments. Upon reaching predetermined locations, it maintained a stationary pose for 60 seconds to facilitate the accumulation of multiple LiDAR scans. The synchronized sensor data, including LiDAR point clouds, camera images, and IMU measurements, were recorded as raw ROS (Robot Operating System) packets and systematically organized into dataset collections. To ensure data diversity and minimize inter-dataset correlation, the robot’s orientation and position were deliberately adjusted between recording sessions, thereby maintaining the independence and representativeness of each dataset instance.
@@ III-A2 Ground Truth Generation
This section exploits two projection methods to generate ground-true LiDAR reflectance image: panoramic spherical projection and virtual-camera perspective projection. In the former one, LiDAR points are directly mapped onto a spherical surface. The azimuth $\theta=\arctan 2(y,x)$ and elevation $\phi=\arctan 2\left(z,\sqrt{x^{2}+y^{2}}\right)$ of a specific 3D point are converted into 2D coordinates: $\text{col}=\left\lfloor\frac{-\theta+180^{\circ}}{360^{\circ}}\cdot W\right\rfloor$ and $\text{row}=\left\lfloor\frac{\phi_{\text{max}}-\phi}{\phi_{\text{max}}-\phi_{\text{min}}}\cdot H\right\rfloor$ where $W$ and $H$ are the width and height of the generated image. $\phi_{\text{min}},\phi_{\text{max}}]$ defines the vertical FoV. The intensity $I$ is assigned to the corresponding pixel $(\text{row},\text{col})$. This method aligns with the LiDAR sensor’s FoV, producing images with a resolution of 1380 × 240 pixels.
In the virtual-camera perspective projection, four virtual cameras are placed around the LiDAR sensor to project the point cloud into four views. Using a pinhole camera model, 3D points $(x,y,z)$ are mapped onto a 2D image plane,
@@
where $f_{x}=\frac{W}{2\cdot\tan\left(\frac{\text{horizontal FoV}}{2}\right)}$, $f_{y}=\frac{H}{\phi_{\text{max}}-\phi_{\text{min}}}$, and $c_{x}=\frac{W}{2}$. Each view generates reflectance images with a resolution of 240 × 480 pixels.
The exemplar reflectance images generated by these two projection modes are shown in Fig. 4. Since the density of LiDAR reflectance image varies with the number of accumulated scans, we use reflectance images accumulated over 500 scans as ground-truth to provide supervision during the training process. Through the aforementioned data acquisition, we constructed the Super LiDAR Reflectance dataset, which includes 1,000 raw ROS packets (600K frames) and an additional 5,000 sets of reflectance images ranging from sparse (1 scan) to dense (500 scans) configuration, covering diverse urban environments such as residential areas, streets, parking lot and intersections. The dataset also includes two test subsets targeting specific applications of dense reflectance images: one is for loop closure detection, with data from three scenarios at two time intervals of day (morning, evening), using panoramic spherical projection for robust matching across different viewpoints. The other subset is designed for lane detection under varying traffic scenarios during the day and night, where the virtual-camera perspective projection is used (more suitable for this task).
Figure 4: Reflectance images obtained by virtual-camera projection mode (top) and panoramic mode (bottom), respectively, when the NRS-LiDAR is placed at a preset location for 60 seconds.

### v2

# Super LiDAR Intensity for Robotic Perception

@@ Abstract
Conventionally, human intuition defines vision as a modality of passive optical sensing, relying on ambient light to perceive the environment. However, active optical sensing, which involves emitting and receiving signals, offers unique advantages by capturing both radiometric and geometric properties of the environment, independent of external illumination conditions. This work focuses on advancing active optical sensing using Light Detection and Ranging (LiDAR), which captures intensity data, enabling the estimation of surface reflectance that remains invariant under varying illumination. Such properties are crucial for robotic perception tasks, including detection, recognition, segmentation, and Simultaneous Localization and Mapping (SLAM). A key challenge with low-cost LiDARs lies in the sparsity of scan data, which limits their broader application. To address this limitation, this work introduces an innovative framework for generating dense LiDAR intensity images from sparse data, leveraging the unique attributes of non-repeating scanning LiDAR (NRS-LiDAR). We tackle critical challenges, including intensity calibration and the transition from static to dynamic scene domains, facilitating the reconstruction of dense intensity images in real-world settings. The key contributions of this work include a comprehensive dataset for LiDAR intensity image densification, a densification network tailored for NRS-LiDAR, and diverse applications such as loop closure and traffic lane detection using the generated dense intensity images. Experimental results validate the efficacy of the proposed approach, which successfully integrates computer vision techniques with LiDAR data processing, enhancing the applicability of low-cost LiDAR systems and establishing a novel paradigm for robotic vision via active optical sensing–LiDAR as a Camera. The dataset and code are available at: https://github.com/IMRL/Super-LiDAR-Intensity.
@@ I Introduction
Recent advancements in LiDAR technology have greatly enhanced its accessibility, driven by reductions in sensor size and cost, thereby expanding its applicability across a wide range of robotic applications. Beyond delivering precise 3D spatial data, LiDAR systems are capable of capturing intensity from objects, acquiring illumination-invariant environmental information, and rich texture to complement the limitations of pure geometric-based LiDAR sensing.
@@
Figure 1: Comparison between the intensity image densified by the proposed method and the true intensity images stitched from high-line LiDAR data. (a) Raw NRS-LiDAR input. (b) Densified 1380 $\times$ 240 intensity image using the proposed method. (c) Stitched 1024 $\times$ 128 intensity images using the high-cost Ouster 128-line LiDAR data from [29, 34]. The proposed method effectively eliminates line artifacts and mitigates overexposure in challenging scenarios, generating high-definition, camera-grade intensity images.
@@
To date, numerous studies have used LiDAR intensity in various applications, e.g., loop closure detection [34, 4], intensity-enhanced LiDAR odometry [29, 53, 6], segmentation [32, 41], and detection [13, 20]. These applications typically depend on dense LiDAR intensity images, which are often generated by high-resolution, costly LiDAR sensors such as Ouster OS1-128 and Velodyne VLS-128. In contrast, low-cost LiDAR sensors, such as Velodyne-16 or Livox MID-360, generally produce sparse data, posing significant challenges for directly applying methods designed for dense intensity images.
@@
Intuitively, generating dense intensity images from sparse LiDAR data holds significant potential for robotic perception, yet this area remains underexplored in existing research [3]. Successfully tackling this challenge would effectively bridge the gap between computer vision and LiDAR sensing techniques, facilitating the adaptation of camera-based methodologies to LiDAR systems. Moreover, it would address the limitations inherent in traditional cameras, such as sensitivity to lighting conditions [11, 12], while simultaneously enhancing the texture representation capabilities of LiDAR technology.
@@
Creating dense LiDAR intensity images typically depends on high-quality, high-resolution ground-truth data, particularly when using deep neural networks. However, acquiring such data often requires specialized sensors or configurations, posing significant challenges. In this study, we capitalize on the distinctive imaging properties of NRS-LiDAR to produce high-quality ground-truth intensity images (Fig. 1). As shown in Fig. 2, dense ground-truth data can be acquired in static environments through scan accumulation. Nevertheless, acquiring dense ground-truth data is not feasible when the robot is in motion. In dynamic scenarios, sparse and rapidly changing inputs must be processed in real time, significantly increasing the complexity of dense reconstruction. The disparity between static ground-truth generation and dynamic scene processing underscores the challenges of achieving dense intensity images with low-cost LiDARs. In summary, we face two key challenges: obtaining dense ground-truth for intensity image densification under motion, and utilizing this ground-truth to densify the sparse intensity image. To tackle these issues, we propose a novel approach with the following key contributions:
@@
    The first benchmark dataset for LiDAR intensity densification. Compared to existing datasets [19, 29], the proposed dataset provides paired sparse-to-dense ground-true data, allowing both learning based densification benchmarking and image-level perception tasks.
@@
    A novel convolutional network for real-time intensity densification for NRS-LiDAR (e.g., Livox MID-360). Our approach effectively addresses two critical challenges: the static-to-motion domain gap and adaptive intensity calibration in LiDAR intensity densification. Compared to existing methods, our method achieves significant improvements in quality and accuracy.
@@
    The validation in loop closure and traffic lane detection tasks. The proposed method is proven effective under challenging conditions, enabling enhanced practical applications for low-cost LiDAR sensors.
@@
Figure 2: Illustration of the non-repeating scanning mechanism. When stationary, the LiDAR gradually produces denser point clouds over time within the same FoV, allowing the intensity image derived from the point cloud projection to achieve higher density as well. The color of point clouds or gray-scale images represents the intensity value.
II Related Work
II-A LiDAR Intensity in Robotics
@@
In single LiDAR-based applications, numerous studies have primarily focused on exploiting geometric information for robotic perception[30] and localization[50]. Nevertheless, these approaches exhibit limitations in geometrically degenerated environments due to their inability to capture texture information. To address this, researchers have explored fusing camera and LiDAR data to exploit their complementary strengths [46, 23]. However, this integration presents challenges, including additional hardware, precise calibration, and temporal synchronization. Moreover, visual sensors are inherently limited by illumination conditions.
@@
A promising alternative focuses on leveraging LiDAR’s intrinsic intensity properties. Significant progress has been demonstrated in this domain. A loop closure detection method that synergistically combines geometric and intensity data is proposed [45]. Furthermore, the application of reflectance information has proven particularly effective in robotic terrain analysis, enabling semantic segmentation to distinguish between road surfaces and grasslands[10]. Although recent learning-based works have advanced intensity simulation by incorporating RGB images and physics-based priors [38, 1], they remain confined to the fidelity of data generation and do not further explore the deployment of such enhanced intensity for downstream robotic tasks.
@@
Recently, dense intensity images, which are generated by high-resolution LiDAR sensors, have been utilized to enhance robotic perception capabilities[53, 6]. A notable work is COIN-LIO[29], which is a LiDAR-inertial odometry framework integrating LiDAR intensity images with geometric registration to address challenges in geometrically degenerated environments. A hybrid loop closure detection method is introduced in [34] by combining LiDAR point cloud and intensity image, achieving superior robustness. However, these advanced methods rely on high-resolution, high-cost LiDAR sensors, making them impractical in resource-constrained scenarios. This study presents a novel approach of reconstructing high-fidelity dense intensity images from sparse ones with low-cost LiDAR, and demonstrates significant potential in real-world real-time robot applications.
II-B Depth and Point Cloud Completion
@@
Dense environmental sensing is critical for robotic tasks. However, physical sensors typically provide only sparse measurements. Completing sparse sensory data into dense perceptual representations is thus a crucial research direction.
Depth completion[37, 27] is one of the most extensively studied domains, inspiring us to review dense completion methods. Depth completion utilizes sparse depth data from sensor measurements[31, 25, 51, 24], multi-view geometry[56, 54], or data-driven priors[49, 28], to generate dense depth images, enabling continuous and dense perception for navigation and manipulation. Beyond depth images, point cloud completion has also garnered significant attention [55]. Recent advances employ attention mechanisms [43], discrete representations [48], or unified diffusion models [5] to reconstruct 3D shapes from partial observations. However, these methods primarily focus on recovering 3D geometric structures, often neglecting the explicit utilization of intensity information which is crucial for visual perception.
@@
Therefore, directly applying existing depth or point cloud completion approaches to LiDAR intensity images densification is challenging due to the complexities of intensity measurements. Factors such as incidence angle, surface properties, and sensor-to-target distance [8] complicate the learning of latent intensity patterns, especially with limited public training data available. Therefore, developing novel networks specifically for LiDAR intensity images densification is essential. Unlike depth completion, which focuses solely on geometry, LiDAR intensity images densification equips robots with a new, high-resolution eye, thus allowing high-level visual perception to be achieved through LiDAR.
@@ III Methodology
Figure 3: System diagram: During stationarity, the robot utilizes NRS-LiDAR to obtain sparse and dense LiDAR data. Sparse data is augmented to simulate motion characteristics, while dense data serves as supervision to train an intensity image densification network for motion perception tasks.
Robots in motion often require sufficient time to accumulate data for a comprehensive understanding of their environment. In this section, we detail our proposed method to address this challenge. As shown in Fig. 3, the purpose of this work is to rely solely on sparse LiDAR data collected by the robot over a short period during motion as input to generate dense and accurate intensity images as output, providing comprehensive environmental information for robotic perception tasks.
@@ III-A1 Data Acquisition
The Super LiDAR Intensity dataset was acquired using a Giraffe ground robot platform equipped with a multi-sensor suite: an Intel RealSense D435i camera capturing RGB images at 1280 × 720 resolution, a Livox MID-360 LiDAR sensor with a 360° horizontal and 59° vertical field of view (FoV) operating at 10 Hz point cloud generation rate, and a high-frequency IMU sampling at 200 Hz. Initially, the robot autonomously navigated through selected representative environments. Upon reaching predetermined locations, it maintained a stationary pose for 60 seconds to facilitate the accumulation of multiple LiDAR scans. The synchronized sensor data, including LiDAR point clouds, camera images, and IMU measurements, were recorded as raw ROS (Robot Operating System) packets and systematically organized into dataset collections. To enhance data diversity and minimize inter-dataset correlation, the robot’s orientation and position were deliberately adjusted between recording sessions. For the same scene, the robot was either moved a distance $d>30\,\mathrm{m}$ or rotated by an angle $\theta>90^{\circ}$ to capture diverse perspectives and ensure sample uniqueness. The dataset comprises 20 distinct urban scenes, including gardens, libraries, plazas, etc., with an average density of approximately 40 samples per kilometer. To eliminate data leakage, the dataset was split by scenes, ensuring that the training, validation, and test sets were collected from entirely separate environments with no overlap.
@@ III-A2 Ground Truth Generation
This section exploits two projection methods to generate ground-true LiDAR intensity image: panoramic spherical projection and virtual-camera perspective projection. In the former one, LiDAR points are directly mapped onto a spherical surface. The azimuth $\theta=\arctan 2(y,x)$ and elevation $\phi=\arctan 2\left(z,\sqrt{x^{2}+y^{2}}\right)$ of a specific 3D point are converted into 2D coordinates: $\text{col}=\left\lfloor\frac{-\theta+180^{\circ}}{360^{\circ}}\cdot W\right\rfloor$ and $\text{row}=\left\lfloor\frac{\phi_{\text{max}}-\phi}{\phi_{\text{max}}-\phi_{\text{min}}}\cdot H\right\rfloor$ where $W$ and $H$ are the width and height of the generated image. $\phi_{\text{min}},\phi_{\text{max}}]$ defines the vertical FoV. The intensity $I$ is assigned to the corresponding pixel $(\text{row},\text{col})$. This method aligns with the LiDAR sensor’s FoV, producing images with a resolution of 1380 × 240 pixels.
In the virtual-camera perspective projection, four virtual cameras are placed around the LiDAR sensor to project the point cloud into four views. Using a pinhole camera model, 3D points $(x,y,z)$ are mapped onto a 2D image plane:
@@
where $f_{x}=\frac{W}{2\cdot\tan\left(\frac{\text{horizontal FoV}}{2}\right)}$, $f_{y}=\frac{H}{\phi_{\text{max}}-\phi_{\text{min}}}$, and $c_{x}=\frac{W}{2}$. Each view generates intensity images in 240 × 480 pixels.
The exemplar intensity images generated by these two projection modes are shown in Fig. 4. Since the density of LiDAR intensity image varies with the number of accumulated scans, we use intensity images accumulated over 500 scans as ground-truth to provide supervision during the training process. Through the aforementioned data acquisition, we constructed the Super LiDAR Intensity dataset, which includes 1,000 raw ROS packets (600K frames) and an additional 5,000 sets of intensity images ranging from sparse (1 scan) to dense (500 scans) configuration, covering diverse urban environments. Additionally, we include specialized subsets: 1) DCM Supervision Data for training the compensation module; 2) Compensation Verification Data; and 3) Application Data targeting specific tasks. The application subset includes data for loop closure detection, with data from three scenarios at two time intervals of the day (morning, evening), using panoramic spherical projection for robust matching across different viewpoints. The other subset is designed for lane detection under varying traffic scenarios during the day and night, where the virtual-camera perspective projection is used (more suitable for this task).
Figure 4: Intensity images obtained by virtual-camera projection mode (top) and panoramic mode (bottom), respectively.